40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7534     Entry time: Fri Oct 12 01:56:26 2012
Author: kiwamu 
Type: Update 
Category: General 
Subject: Alignment situation of interferometer  

[Koji / Kiwamu]

 We have realigned the interferometer except the incident beam.

 The REFL beam is not coming out from the chamber and is likely hitting the holder of a mirror in the OMC chamber.

So we need to open the chamber again before trying to lock the recycled interferometers at some point.

 

--- What we did

  •  Ran the MC decenter script to check the spot positions.
    • MC3 YAW gave a - 5mm offset with an error of about the same level.
    • We didn't believe in this dither measurement.
  •  Checked the IP-POS and IP-ANG trends.
    • The trends looked stable over 10 days (with a 24 hours drift).
    • So we decided not to touch the MC suspensions.
  • Tried aligning PRM
  • Found that the beam on the REFL path was a fake beam
    • The position of this beam was not sensitive to the alignment of PRM or ITMs.
    • So certainly this is not the REFL beam.
    • The power of this unknown beam is about 7.8 mW
  • Let the PRM reflection beam go through the Faraday
    • This was done by looking at the hole of the Faraday though a view port of the IOO chamber with an IR viewer.
  • Aligned the rest of the interferometer (not including ETMs)
    • We used the aligned PRM as the alignment reference
    • Aligned ITMY such that the ITMY reflection overlaps with the PRM beam at the AS port.
    • Aligned the BS and SRM such that their associated beam overlap at the AS port
    • Aligned ITMX in the same way.
    • Note that the beam axis, defined by the BS, ITMX  and SRM, was not determined by this process. So we need to align it using the y-arm as a reference at some point.
    • After the alignment, the beam at the AS port still doesn't look clipped. Which is good.

 

---- things to be fixed

   - Align the steering mirrors in the faraday rejected beam path (requires vent)

   - SRM oplev (this is out of the QPD range)

   - ITMX oplev (out of the range too)

ELOG V3.1.3-